OpenStudy (sasogeek):

find the coefficient of \(\large x^4 \) in the binomial of \(\large (3-2x)^6 \)

OpenStudy (mimi_x3):

Is in the expansion and binomial the same? lol

OpenStudy (sasogeek):

i believe so

OpenStudy (sasogeek):

here's what i'm getting though it's not in the possible answers...270 that being \[\ 15(3^2)\times 2 \]

OpenStudy (anonymous):

let evaluate this way: we know (n,k)=nCk

OpenStudy (sasogeek):

oh i see my mistake... i forgot to raise 2 to the power 4 too :) let's see if it get it :)

OpenStudy (mimi_x3):

If yes, then.. \[\left(\begin{matrix}6 \\ 4\end{matrix}\right)*(3)^{6-4}*(-2x)^4\] \[15*3^2*16x^4\] = \[2160x^4\]

OpenStudy (sasogeek):

yea 2160 is a possible solution :)

OpenStudy (mimi_x3):

Yes its correct. http://www.wolframalpha.com/input/?i=%283-2x%29^6

OpenStudy (anonymous):

also we know (a+b)^2=2C0a^2+2C1ab+2C2b^2 so do the same instead of power six go up to coef of 4

OpenStudy (anonymous):

this will be good web to look this qn. well documented http://www.themathpage.com/aprecalc/binomial-theorem.htm

OpenStudy (mimi_x3):

Well, you can use the formula, that i used that is easier i guess..